You have a jar containing 36 coins, all of which are quarters and nickels. The total value is $4.40. How many of each type of coin is there?
Answer by CubeyThePenguin(3113)

q = number of quarters
n = number of nickels

q + n = 36
25q + 5n = 440

Multiply the first equation by 25.

25q + 25n = 900
-(25q + 5n = 440)
----------------------
20n = 460
n = 23

There are 23 nickels and 13 quarters.